OpenStudy (anonymous):

Please help!!! I'll give medals! Max has just won some money on a game show! He has the option to take a lump sum payment of $500,000 now or get paid an annuity of $4,900 per month for the next 10 years. Assuming the growth rate of the economy is 2.9% compounding annually over the next 10 years, which is the better deal for Max and by how much? Lump Sum: by $77,462.75 Lump Sum: by $4,145.41 Annuity: by $88,000.00 Annuity: by $4,145.41

OpenStudy (anonymous):

$500,000 = lump of sum $4,900 * 12 = 58,800 $58,800 * 10 = 588,000 (if we leave it like this the answer would be C;Annuity: by $88,000.00, but what i think we also have to add is 2.9 * 10 = 29 {since the ecnomic growth is 2.9 a year} and times that but 588,000 which equal out to be 17, 052, 000 which doesn't get answer... or just times 588,000 by 2.9 which equals 1,702,000 and doesn't get an answer... maybe i'm thinking to much...sorry if this didn't help :( )

OpenStudy (anonymous):

Thank you for trying anyway! I'm going to put C because I can't get it to make sense either.

OpenStudy (anonymous):

welcome! and sorry! ^^